OpenStudy (anonymous):

Find the exact value by using a half-angle identity. tangent of seven pi divided by eight.

OpenStudy (mathstudent55):

\(\large \tan \dfrac{7 \pi}{8} = \tan \dfrac{\frac{7 \pi}{4}}{2}\)

OpenStudy (mathstudent55):

\(\tan \dfrac{\theta}{2} = \dfrac{1 - \cos \theta}{\sin \theta} \) \(\cos \dfrac{7\pi}{4} \) and \(\sin \dfrac{7\pi}{4} \) are well known.

OpenStudy (anonymous):

1-sqrt2